Please solve the following system of linear equations with substitution.
3x-2=y
y-2x=-5

Answer by ankor@dixie-net.com(22740)

3x - 2 = y
y - 2x =-5
From the 1st equation we know that y = 3x - 2, therefore we can replace y with (3x-2) in the 2nd equation
(3x-2) - 2x = -5
3x - 2x = -5 + 2
x = -3
:
Find y
y = 3x - 2
we know x=-3
y = 3(-3) - 2
y = -9 - 2
y = -11
:
:
Check our solutions of x=-3, y=-11 in the 2nd equation
y - 2x = -5
replace x with -3, y with -11
-11 - 2(-3) = -5
minus a minus is a plus so we have
-11 + 6 = -5; which is true so our solution check out